ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

5. Differential Geometry and Global Analysis

1991 MS Classification: 053 Yevtouchik, Leonid, Moscow State University, Russia Nonlinear connections and its applications Nonlinear stabile connections γ^{p} introduced by the author in higher order frames bundles H^{p}(X_{n}) induce on X_{n} a system of ordinary differential equations of the order p + 1:

S^{p+1} : T^{p}(X_{n}) → T^{p+1}(X_{n}) .

A natural differential operator from the set of arbitrary stationar systems S^{p+1} in the set of stabile nonlinear connections γ^{p} is constructed.